To furnish the teacher’s cafeteria in a school the director ordered some chairs for 40,000 L.L each ,and some tables for 16,000 L.L each. he paid a total of 1,880,000L.L how many chairs and how many tables did he order if the number of chairs exceeds the number of tables by five?
Answer by mananth(16946)

chairs for 40,000 L.L each ,and some tables for 16,000 L.L each. he paid a total of 1,880,000L.L how many chairs and how many tables did he order if the number of chairs exceeds the number of tables by five?
let number of tables bought be x
number of chairs is 5 more so (x+5)
16000x+40000(x+5)=1,880,000
16000x+40000x+200000=1880000
56000x =1880000-200000
56000x=1680000
x=30
plug x to get y
y=35